If the sides of a square are decreased by 2cm, the area is decreased by 36cm^2. What are the dimensions of the original square?

Answer by stanbon(75887)

If the sides of a square are decreased by 2cm, the area is decreased by 36cm^2. What are the dimensions of the original square?
------------------
Let the original square have sides that are "x"cm. long.
The area of this square is x^2 cm^2
----------------
The new sqquare has sides that are "x-2"cm. long.
The area of this square is (x-2)^2 = x^2-4x+4 cm^2
------------
EQUATION:
original area minus new area = 36 cm^2
x^2-(x^2-4x+4) = 36
4x-4 = 36
4x = 40
x = 10 cm
The original square is 10 cm by 10 cm
=====================
Cheers,
Stan H.